OpenStudy (anonymous):

what is the solution of 7k+2=4k-10? A)k=-12 B)k=-4 C)k=-11/12 D)k=4

OpenStudy (anonymous):

it's b k=-4 i checked

OpenStudy (anonymous):

so did you get it right

OpenStudy (swamijatin2):

7k +2=4k-10 7k-4k=-2-10 3k=-12 k=-4

OpenStudy (anonymous):

i'm write and ha Dexster hows your sister still driving you crazy